# Quantum Blackwell’s Ordering and Differential Privacy

### v1

@@ VI Quantum Privatized Parameter Inference
More generally, in quantum parameter estimation, the goal is to estimate an unknown parameter $\theta$ encoded in a family of quantum states $\{\rho_{\theta}\}$, where the encoding is constrained by privacy requirements. In the privatized setting, the states $\rho$ and $\sigma$ must satisfy the $(\varepsilon,\delta)$-DP condition, which restricts the distinguishability of the states and thus limits the amount of information that can be extracted about $\theta$.
The estimation performance is fundamentally limited by the quantum Fisher information, which quantifies the sensitivity of the state $\rho_{\theta}$ to changes in $\theta$. In the quantum setting, the relevant quantity is the Symmetric Logarithmic Derivative (SLD) Fisher information, defined as
$J_{\theta}:=\mathrm{Tr}\left[L_{\theta}^{2}\rho_{\theta}\right],$
where $L_{\theta}$ is the SLD operator satisfying
$\frac{1}{2}(L_{\theta}\rho_{\theta}+\rho_{\theta}L_{\theta})=\frac{d\rho_{\theta}}{d\theta}.$
The quantum Cramér-Rao bound states that the mean squared error (MSE) of any unbiased estimator $\hat{\theta}$ is lower bounded by $1/J_{\theta}$.
In the privatized scenario, the investigator is restricted to the pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$, and thus the achievable Fisher information is maximized over all such admissible pairs. This leads to a constrained optimization problem,
$\max_{(\rho,\sigma):(\varepsilon,\delta)\text{-DP}}J_{\theta},$
@@
The investigator’s goal is to estimate the parameter $\theta$. The quality of this estimation is typically measured by the mean square error (MSE). For a single-parameter model, the MSE is lower-bounded by the Cramér-Rao bound, which is the inverse of the Fisher information. In the quantum setting, we use the Symmetric Logarithmic Derivative (SLD) Fisher information. The SLD, denoted $L_{\theta}$, is defined by the equation:
@@
The MSE of any unbiased estimator for $\theta$ is lower-bounded by $1/J_{\theta}$. This bound is asymptotically achievable, for instance, via a two-step estimation process. Consequently, to optimize the estimation, it is natural to maximize the SLD Fisher information $J_{\theta}$ subject to the $(\varepsilon,\delta)$-DP privacy constraint on the states $\rho$ and $\sigma$.
@@
$\max_{(\rho,\sigma):(\varepsilon,\delta)\text{-DP}}J_{\theta}=\frac{\delta}{\theta(1-\theta)}+\frac{(1-\delta)(1-e^{\varepsilon})^{2}}{e^{\varepsilon}+(1-e^{\varepsilon})^{2}\theta(1-\theta)}.$ (68)
@@
Let $\rho_{(\varepsilon,\delta),\theta}:=\theta\rho_{(\varepsilon,\delta)}+(1-\theta)\sigma_{(\varepsilon,\delta)}$. We first compute the SLD Fisher information for this specific state, which we denote by $J_{(\varepsilon,\delta),\theta}$. The state $\rho_{(\varepsilon,\delta),\theta}$ is diagonal in the standard basis $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$. Its eigenvalues are:

### v3

@@ V-A Problem setting
More generally, quantum parameter inference concerns estimating an unknown parameter $\theta$ encoded in a family of states $\{\rho_{\theta}\}$, where the encoding is constrained by privacy requirements. In the present privatized setting, the requirement $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ restricts the distinguishability of the two single-shot outputs under any measurement and hence limits how much information about $\theta$ can be extracted from $\rho_{\theta}^{\otimes n}$. In the following subsections, we quantify this limitation through (i) privatized hypothesis testing and (ii) bounds on the (SLD) Fisher information under $(\varepsilon,\delta)$-QDP.
V-B Hypothesis testing with one respondent
First, for the theoretical and fundamental analysis, we consider the scenario where the investigator aims to distinguish between two original information sources, $\rho_{\theta_{0}}$ and $\rho_{\theta_{1}}$ with $\theta_{0}<\theta_{1}$, only from one observation; however, this scenario is impractical as it addresses only a single respondent. The performance in this hypothesis testing task is characterized by quantities like the hypothesis testing divergence $D_{H}^{\alpha}(\rho_{\theta_{0}}\|\rho_{\theta_{1}})$.
Theorem 3.
$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}D_{H}^{\alpha}(\rho_{\theta_{1}}\|\rho_{\theta_{2}})=D_{H}^{\alpha}(\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}}).$ (77)
Proof.
It is sufficient to show
$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}\beta^{c}(\alpha|\rho_{\theta_{1}}\|\rho_{\theta_{0}})=\beta^{c}(\alpha|\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}}).$ (78)
where $\beta^{c}(\alpha|\sigma_{1}\|\sigma_{2}):=1-\beta(\alpha|\sigma_{1}\|\sigma_{2})$. The relation (61) implies
$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}\beta^{c}(\alpha|\rho\|\sigma)=\beta^{c}(\alpha|\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)}).$ (79)
Hence, the following discussion will show (78) from (79).
Assume that the states $\rho$ and $\sigma$ satisfy the $(\varepsilon,\delta)$-DP condition, which implies
$\beta^{c}(\alpha|\rho\|\sigma)\leq\beta^{c}(\alpha|\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)}).$ (80)
The quantity $\beta^{c}(\alpha|\rho_{\theta_{1}}\|\rho_{\theta_{0}})$ is written as
$\beta^{c}(\alpha|\rho_{\theta_{1}}\|\rho_{\theta_{0}})=$ $\max_{\begin{subarray}{c}\alpha^{\prime}\in[0,1]:\\ \theta_{1}\alpha^{\prime}+(1-\theta_{1})\beta^{c}(\alpha^{\prime}|\rho\|\sigma)\leq\alpha\end{subarray}}\theta_{0}\alpha^{\prime}+(1-\theta_{0})\beta^{c}(\alpha^{\prime}|\rho\|\sigma).$ (81)
We choose a real number $\alpha_{0}$ satisfying $\theta_{1}\alpha_{0}+(1-\theta_{1})\beta^{c}(\alpha_{0}|\rho\|\sigma)=\alpha$. Then, $\beta^{c}(\alpha|\rho_{\theta_{1}}\|\rho_{\theta_{0}})=\theta_{0}\alpha_{0}+(1-\theta_{0})\beta^{c}(\alpha_{0}|\rho\|\sigma)$ because $\theta_{0}\alpha^{\prime}+(1-\theta_{0})\beta^{c}(\alpha^{\prime}|\rho\|\sigma)$ and $\theta_{1}\alpha^{\prime}+(1-\theta_{1})\beta^{c}(\alpha^{\prime}|\rho\|\sigma)$ are strictly monotonically increasing for $\alpha^{\prime}$. Also, we choose a real number $\alpha_{1}$ satisfying $\theta_{1}\alpha_{1}+(1-\theta_{1})\beta^{c}(\alpha_{1}|\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})=\alpha$. Hence, the relation (80) implies that $\alpha_{1}\leq\alpha_{0}$. Hence,
$\beta^{c}(\alpha|\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}})-\beta^{c}(\alpha|\rho_{\theta_{1}}\|\rho_{\theta_{0}})$
$=$ $\theta_{0}\alpha_{1}+(1-\theta_{0})\beta^{c}(\alpha_{1}|\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}})-(\theta_{0}\alpha_{0}+(1-\theta_{0})\beta^{c}(\alpha_{0}|\rho\|\sigma))$
$=$ $(\theta_{1}-\theta_{0})(\beta^{c}(\alpha_{1}|\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}})-\alpha_{1})-(\theta_{1}-\theta_{0})(\beta^{c}(\alpha_{0}|\rho\|\sigma)-\alpha_{0})$
$=$ $(\theta_{1}-\theta_{0})(\beta^{c}(\alpha_{1}|\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}})-\beta^{c}(\alpha_{0}|\rho\|\sigma)+\alpha_{0}-\alpha_{1})\geq 0,$
which implies (78). This completes the proof of Theorem 3.
The combination of Theorems 1 and 3 derives the following corollary.
Corollary 4.
For any $\alpha\in[0,1)\cup(1,+\infty)$, we have
$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}D_{\alpha}(\rho_{\theta_{1}}\|\rho_{\theta_{0}})$ $=D_{\alpha}(\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}}),$ (82)
$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}D(\rho_{\theta_{1}}\|\rho_{\theta_{0}})$ $=D(\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}}),$ (83)
$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}\check{D}_{\alpha}(\rho_{\theta_{1}}\|\rho_{\theta_{0}})$ $=\check{D}_{\alpha}(\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}}),$ (84)
V-C Hypothesis testing with $n$ respondents
Next, we discuss the hypothesis testing when the investigator receives $n$ responses. This scenario focuses on the hypothesis testing divergence $D_{H}^{\alpha}(\rho_{\theta_{1}}^{\otimes n}\|\rho_{\theta_{0}}^{\otimes n})$. We study the following quantity
$B_{\epsilon}(\rho_{\theta_{1}}\|\rho_{\theta_{0}})$ $:=\lim_{n\to\infty}\frac{1}{n}D_{H}^{\epsilon}(\rho_{\theta_{1}}^{\otimes n}\|\rho_{\theta_{0}}^{\otimes n}),$ (85)
for $\epsilon\in(0,1)$. Since it is known that the above values are given by $D(\rho_{\theta_{1}}\|\rho_{\theta_{0}})$ as Stein’s lemma [17, 27], [16, Chapter 3], thus, (83) of Corollary 4 yields the following relation.
$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}B_{\epsilon}(\rho_{\theta_{1}}\|\rho_{\theta_{0}})$ $=D(\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}}).$ (86)
The above relations show that the asymptotic behavior of the hypothesis testing problem is optimized when respondents encode their information by using $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$.
V-D Parameter estimation
@@
Next, we discuss the estimation of the parameter $\theta$ when the investigator has $n$ data. The investigator’s goal is to estimate the parameter $\theta$. The quality of this estimation is typically measured by the mean square error (MSE). For a single-parameter model, the MSE is lower-bounded by the Cramér-Rao bound, which is the inverse of the Fisher information. In the quantum setting, we use the Symmetric Logarithmic Derivative (SLD) Fisher information. The SLD, denoted $L_{\theta}$, is defined by the equation:
@@
The MSE of any unbiased estimator for $\theta$ is lower-bounded by $1/J_{\theta}$. This bound is asymptotically achievable, for instance, via a two-step estimation process.
In the privatized scenario, the investigator is restricted to the pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$. Consequently, to optimize the estimation, it is natural to maximize the SLD Fisher information $J_{\theta}$ subject to the $(\varepsilon,\delta)$-DP privacy constraint on the states $\rho$ and $\sigma$. This leads to a constrained optimization problem,
$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}J_{\theta},$
@@
$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}J_{\theta}=\frac{\delta}{\theta(1-\theta)}+\frac{(1-\delta)(1-e^{\varepsilon})^{2}}{e^{\varepsilon}+(1-e^{\varepsilon})^{2}\theta(1-\theta)}.$ (89)
@@
Let $\rho_{(\varepsilon,\delta),\theta}:=\theta\rho_{(\varepsilon,\delta)}+(1-\theta)\sigma_{(\varepsilon,\delta)}$. First, we denote the SLD Fisher information for this specific states $\rho_{(\varepsilon,\delta),\theta}$ by $J_{(\varepsilon,\delta),\theta}$, and we will show the relation
$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}J_{\theta}=J_{(\varepsilon,\delta),\theta}.$ (90)
It is known that [16, Theorem 6.3]
$J_{\theta}=\lim_{\epsilon\to 0}\frac{8(1-\mathrm{Tr}[|\sqrt{\rho_{\theta}}\sqrt{\rho_{\theta+\epsilon}}|])}{\epsilon^{2}}.$ (91)
Since $\check{D}_{\frac{1}{2}}(\rho\|\sigma)=-\frac{1}{2}\log\mathrm{Tr}[|\sqrt{\rho}\sqrt{\sigma}|]$, the above relation is rewritten as
$J_{\theta}=\lim_{\epsilon\to 0}\frac{4\check{D}_{\frac{1}{2}}(\rho_{\theta}\|\rho_{\theta+\epsilon})}{\epsilon^{2}}.$ (92)
Therefore, the relation (84) with the limit like (92) implies (90).
Next, we compute the SLD Fisher information $J_{(\varepsilon,\delta),\theta}$. The state $\rho_{(\varepsilon,\delta),\theta}$ is diagonal in the standard basis $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$. Its eigenvalues are: